# FIHA: Autonomous Hallucination Evaluation in Vision-Language Models with Davidson Scene Graphs

### v1

@@ 3.1 Setup
Datasets. We construct a hallucination evaluation benchmark FIHA-v1 based on three datasets: the MSCOCO [27], the Foggy [7] and Visual Genome [22]. MSCOCO is a large image dataset developed by Microsoft, officially known as Microsoft Common Objects in Context. This dataset aims to advance the development of computer vision tasks such as object detection, segmentation, and image captioning. This dataset contains over 330,000 images, of which more than 200,000 images are annotated, covering 80 different object categories. Foggy is a synthetic fog dataset that simulates fog in real-world scenes. Each foggy image is rendered using clear images and depth maps from Cityscapes. There are a total of 1,500 images and 500 for each three different foggy levels of each cityscape image: no foggy, medium foggy, and dense foggy. Visual Genome [22] is a large image datasets which includes 108,077 images. Those images have some overlap with the MSCOCO.
Metrics. We use Accuracy (Acc.), Precision (P.), Recall (R.), and F1 Score (F1) as evaluation metrics for Yes-No questions. For Wh-Questions, we use the Accuracy, Recall, and F1 Score from BERTScore [43] for evaluation.
Models We select seven mainstream LVLMs for evaluation: mPLUG-Owl [40], MiniGPT-4 [45], MultiModal-GPT [12], LLaVA-1.5-7B [30], LLaVA-1.5-13B [30], InstructBLIP [10], and GPT-4V [32].
@@ 3.2 Data Processing and Analysis
Figure 3: Distribution of two types of question, i.e. Yes-No and Wh-questions.
@@
Table 4: The results of a more fine-grained assessment of LVLMs from the perspectives of object, attribute, and relation. The results are based on statistics from Q&A pairs generated from captions in the MSCOCO and Foggy Cityscapes dataset.
@@ 3.3.2 Fine-Grained Results
Furthermore, we evaluate the model’s performance from more dimensions (i.e. the object existence, attribute, and relation) with FIHA. We show the fine-grained evaluation results in Table 4.
@@ 4.1 How Reliable is FIHA?
To assess the reliability of the benchmark, we manually evaluate the accuracy of the Q&A pairs in the FIHA-v1 generated by the pipeline. This evaluation involves verifying the correctness of the Q&A pairs by checking whether the answers match the corresponding questions. We show the evaluation results in Table 5, showing that the Q&A pairs generated from image captions are highly reliable, achieving a 96% accuracy rate in samples from the MSCOCO dataset. The pipeline based on MSCOCO, which uses Fast R-CNN, achieves a 98.2% accuracy rate for Q&A pairs. However, the remaining 1.8% error can lead to incorrect Q&A pairs. For example, the model sometimes fails to identify specific details, such as ears in an image. Another challenge relates to feature extraction. For instance, in one case, a horse’s color is incorrectly identified as black when it is actually white or light gray. Although such inaccuracies are rare, they highlight a weakness of the framework and underscore the need for ongoing advancements in object detection and feature extraction technologies. Overall, the FIHA framework demonstrates high reliability in generating datasets for evaluating hallucinations in LVLMs, with the dataset generated from captions performing exceptionally well and achieving near-perfect accuracy.
Table 5: The accuracy of Q&A pairs generated from different datasets.
@@
In addition, we conduct tests on more complicated images. Specifically, we use the Foggy dataset [7], which contains three-level foggy for an image: no foggy, medium foggy, and dense foggy. We show the cases in this dataset in Appendix B. We select a total of 150 images, with 50 images for each of the three foggy levels. This comparison aims to assess the influence of noise on the accuracy of the framework. For the Q&A pair generation from images approach, as shown in Table 5, in the dense fog condition, the accuracy for Q&A pairs generation from images approach is as high as 96.3%. For Q&A pairs generation from captions approach, the accuracy is 94.1%. For medium fog conditions, it is 97.6% and 94.5%; and for no fog conditions, the accuracy are 98.1% and 94.1%. The no-fog result show a similar accuracy with MSCOCO. From the results, we can see that as the degree of fog increases, the accuracy of FIHA’s Q&A pairs decreases, indicating the challenge posed by blurry images. Notably, despite the image with dense fog, Q&A pairs based on our FIHA still has a high accuracy rate.
4.2 What is the Impact of Introducing the DSG?
To reasonably improve the reliability of hallucination assessment, we introduce the DSG mechanism, which can model the dependency between different types of hallucination. As introduced in the previous section, by reorganizing the problem into a tree structure, the judgment of each leaf node depends on the correctness of the root node’s judgment. In this section, we quantitatively analyze the impact brought by the DSG.
Table 6 shows the performance difference for each model before and after introducing DSG. It shows that stronger models like GPT-4V and LLaVA are less affected, while the metrics for other models have dropped by more than 50%. The reason might be that these weaker models make a lot of mistakes in the root-level questions which mostly ask about the object’s existence and answer the leaf-level questions correctly by random guessing. Therefore, after the introduction of the DSG, they are marked as failing on all related leaf node questions, leading to a significant drop in performance. This indicates the reliability of introducing the DSG mechanism.
@@ 4.2 What is the Impact of Introducing the DSG?
Model Acc.$\downarrow$ P.$\downarrow$ R.$\downarrow$ F1$\downarrow$
mPLUG-Owl 29.6 22.1 14.0 28.7
MiniGPT-4 62.6 51.8 62.1 61.2
MultiModal-GPT 21.3 27.6 21.9 24.3
LLaVA-1.5-7B 4.2 11.7 4.5 4.8
LLaVA-1.5-13B 2.7 8.1 3.3 3.6
InstructBLIP 5.7 9.6 5.7 5.7
GPT-4V 6.0 9.9 5.4 8.4
@@ 4.4 Why are Performance on our Benchmark Lower Than Others?
In the experiment, we observe that our test results are lower than others, e.g. POPE [26] and HaELM [38], indicating that FIHA can detect more difficult and distinct issues. We analyze that there are mainly three reasons: firstly, we added a large number of misleading negative samples, and since the model tends to give affirmative answers (Section 3.3.1), this increases the difficulty of evaluation. Secondly, the role of DSG directly impacts the results and improves the reliability of the evaluation method. Finally, the comprehensiveness of FIHA is more challenging than methods that focus primarily on generating coarse-grained object-level questions.
@@ 5 Related Work
5.1 Large Vision-Language Model
@@
With the success of pretraining techniques in Large Language Models (LLMs) [36] and Vision Foundation Models (VFMs) [3], many researchers [2, 23] expanded LLMs to comprehend real-world images through LVLMs with advance in-context or few-shot learning capabilities. As a result, there has been a surge in visual instruction-adapted LVLMs [30, 45, 10, 12], demonstrating remarkable generalization performance across various Vision-Language (VL) tasks. Most of these studies utilized GPT-4 to generate multimodal instruction tuning datasets and multi-stage pretraining to align the visual information with the pre-trained LLM. For example, Liu et al. [30] utilized the visual encoder output as input for LLaMA [36] and trained both networks to align on the generated visual instruction dataset. Zhu et al. [45] integrated Vicuna [33] as a language decoder and only fine-tuned the cross-modal alignment network with extended image captions from ChatGPT. Likewise, both Multimodal GPT [12] and InstructBLIP [10] used various instruction-tailored VL datasets. In addition, the former adopted BLIP2 [25] as its foundational architecture while the latter initialized from Flamingo [2].
Despite the advancements of LVLMs, they remain encumbered by the persistent challenge of hallucinations when generating textual output. These issues significantly hinder their effectiveness in various vision-language tasks [35].
5.2 Hallucination in LVLMs
Recently, there has been growing research attention directed towards the phenomenon of hallucination in LVLMs. Among these works, some studies, as shown in Table 1, have concentrated on hallucination detection and evaluation [26, 38, 37, 20], and some have developed methods to mitigate hallucination [29, 44, 41, 19]. For example, the Bingo [9] benchmark evaluates hallucinations in GPT-4V by focusing on bias and interference, such as regional bias and vulnerability to leading questions. HallusionBench [13] provides a comprehensive diagnostic suite for evaluating entangled language hallucination and visual illusions, emphasizing visual commonsense and reasoning with diverse visual-question control pairs. AutoHallusion [39] introduces an automatic method for generating hallucination benchmarks by manipulating images to conflict with language priors, evaluating models on both synthetic and real-world data. Hal-Eval [18] offers a universal framework categorizing hallucinations into objects, attributes, relations, and events, utilizing a detailed taxonomy to assess LVLMs. Though the issue of hallucination is studied extensively, only a few works have focused on fine-grained hallucination detection in LVLMs. For instance, Li et al. [26] proposed a novel evaluation metric “POPE” to evaluate hallucinations in LVLMs by pooling questions about the ground truth objects. They showed that existing state-of-the-art LVLMs are highly prone to object-level hallucinations. Wang et al. [38] introduced “HaELM”, a framework for detecting hallucinations. They utilized LLMs to generate a hallucinatory dataset and then fine-tuned LLaMA to identify hallucinatory responses from LVLMs. The aforementioned line of research either exclusively focused on object-level hallucination or required training for the detection of hallucination. To address these challenges, Wang et al. [37] introduced “AMBER”, a comprehensive benchmark capable of assessing both generative and discriminative tasks based on object attribute and relation hallucination. Though this work developed a fine-grained hallucination evaluation framework, it required human annotators to annotate the object existence, object attribute, and object relation information for discriminative tasks.
In contrast to the aforementioned studies, our work differs by being applicable to any existing dataset or unseen images for generating probing questions related to object existence, attributes, and relations for evaluating LVLM hallucination. Instead of relying on human annotators, we use an object detection model that performs better in object detection tasks than LLMs. In addition, we consider the dependent relation between object, attribute, and relation.

### v2

@@ 3.1 Setup
Datasets. We construct a hallucination evaluation benchmark FIHA-v1 based on three datasets: the MSCOCO Lin et al. (2014), the Foggy Cordts et al. (2016) and Visual Genome Krishna et al. (2017). MSCOCO is a large image dataset by Microsoft with over 330,000 images. More than 200,000 are annotated across 80 object categories. Foggy is a synthetic fog dataset with 1,500 images, each in three fog levels (no fog, medium fog, dense fog). Visual Genome has 108,077 images with some overlap with MSCOCO. For our benchmark, we only use the test sets of these datasets to avoid overlap with training data used by LVLMs.
Metrics. We use Accuracy (Acc.), Precision (P.), Recall (R.), and F1 Score (F1) as evaluation metrics for Yes-No questions. For Wh-Questions, we use the F1 (Gen) from BERTScore Zhang et al. (2020) for evaluation.
Models. We select seven mainstream LVLMs for evaluation: mPLUG-Owl Ye et al. (2023), MiniGPT-4 Zhu et al. (2023), MultiModal-GPT Gong et al. (2023), LLaVA-1.5-7B Liu et al. (2023c), LLaVA-1.5-13B Liu et al. (2023c), InstructBLIP Dai et al. (2023), and GPT-4V OpenAI (2024).
@@ 3.2 Data Processing and Analysis
Figure 3: Distribution of two types of question, i.e., Yes-No and Wh-questions.
@@
Table 4: The fine-grained assessment of LVLMs evaluates object, attribute, and relation accuracy using Q&A pairs generated from captions in the MSCOCO and Foggy Cityscapes datasets.
@@ 3.3.2 Fine-Grained Results
Furthermore, we evaluate the model’s performance from more dimensions (i.e., the object existence, attribute, and relation) with FIHA. We show the fine-grained evaluation results in Table 4.
@@ 4.1 How Reliable is FIHA?
To evaluate the benchmark’s reliability, we manually check the accuracy of Q&A pairs in FIHA-v1 generated by the pipeline, verifying if the answers match the questions. In the human evaluation process, we employ annotators manually evaluate whether the answer is right in each image, question, and answer pair. Whether it is a Yes-No question or a Wh-Question, it will be marked as True or False, that is, whether the answer is right for the question. The final accuracy is calculated as (num_true / (num_true + num_false)). Table 5 shows that Q&A pairs from image captions are 96% accurate on MSCOCO samples. The MSCOCO-based pipeline using Grounding DINO achieves 98.2% accuracy, with 1.8% errors, such as missing details or misidentifying colors. Overall, FIHA shows high reliability in generating datasets for evaluating hallucinations in LVLMs, with near-perfect accuracy in caption-based datasets.
Table 5: The results of human evaluation (accuracy) of Q&A pairs generated from different datasets.
@@
We also test on complex images using the Foggy dataset Cordts et al. (2016), which test evaluates noise’s effect on the framework’s accuracy. Examples are in Appendix B.
As shown in Table 5, under dense fog, the accuracy for Q&A pairs generation from images is 96.3%, while for Q&A pairs generation from captions, it is 94.1%. For medium fog, the accuracies are 97.6% and 94.5%, respectively. Under no fog, the accuracies are 98.1% and 94.1%, similar to MSCOCO results.
The results show that as fog increases, the accuracy of FIHA’s Q&A pairs decreases, highlighting the challenge of blurry images.
4.2 What is the Impact of Introducing DSG?
To improve hallucination assessment, we propose the DSG mechanism, which models dependencies between hallucination types. Table 6 shows that stronger models like GPT-4V and LLaVA-1.5-13B exhibit smaller performance drops (6.0% and 2.7%), indicating their robustness to dependency-based evaluations. In contrast, weaker models such as MultiModal-GPT and mPLUG-Owl show substantial declines (21.3% and 29.6%), reflecting frequent root-level errors that propagate to leaf-level questions, which these models might otherwise answer correctly under standard evaluation.
The results reveal that most hallucination errors occur at fundamental levels, such as object recognition, with models like LLaVA-1.5-13B maintaining high accuracy due to fewer cascading errors. Precision drops more than recall across models, suggesting that DSG effectively uncovers false positives. For example, MiniGPT-4’s precision decreased by 11.0%, highlighting previously unnoticed errors.
While InstructBLIP shows a significant F1 drop (from 9.6% to 5.7%), GPT-4V remains relatively stable (9.9% to 8.4%), suggesting stronger contextual reasoning. These results demonstrate that DSG provides a rigorous evaluation, exposing model weaknesses that standard assessments may miss.
@@ 4.2 What is the Impact of Introducing DSG?
Model Acc.$\downarrow$ P.$\downarrow$ R.$\downarrow$ F1$\downarrow$ F1 (Gen)$\downarrow$
mPLUG-Owl 29.6 22.1 14.0 28.7 14.2
MiniGPT-4 62.6 51.8 62.1 61.2 42.3
MultiModal-GPT 21.3 27.6 21.9 24.3 12.9
LLaVA-1.5-7B 4.2 11.7 4.5 4.8 5.7
LLaVA-1.5-13B 2.7 8.1 3.3 3.6 5.1
InstructBLIP 5.7 9.6 5.7 5.7 6.9
GPT-4V 6.0 9.9 5.4 8.4 3.9
4.3 How and Why LLM is Free?
FIHA has a big benefit: it doesn’t need extra big language models like GPT-4. This is shown when we make questions from true information. We do this with Python code, and you can find more details in Section 2.2 and Appendix A.
We don’t use big language models to help make questions because they cost a lot. For example, if someone wants to make questions for 500 pictures, they would need about 36,900 questions (like we saw with MS COCO). This would cost almost \$400 in API fees, which is very expensive.
@@ 4.5 Why are Performance on our Benchmark Lower Than Others?
In the experiment, we observe that our test results are lower than others, e.g., POPE Li et al. (2023d) and HaELM Wang et al. (2023b), indicating that FIHA can detect more difficult and distinct issues. We analyze that there are mainly three reasons: firstly, we added a large number of misleading negative samples, and since the model tends to give affirmative answers (Section 3.3.1), this increases the difficulty of evaluation. Secondly, the role of DSG directly impacts the results and improves the reliability of the evaluation method. Finally, the comprehensiveness of FIHA is more challenging than methods that focus primarily on generating coarse-grained object-level questions.
4.6 Will Using Fixed Templates Limit the Diversity and Types of Questions?
According to the description of LVLM hallucinations in the existing work Bai et al. (2024), the types of questions usually limited, even if the questions are generated by LLMs. We compared the method Chen et al. (2024); Jiang et al. (2024b) of using LLMs to generate questions and found that the diversity of questions generated by the LLM-based method is similar to our method.
4.7 Why Are the Metrics Lower on the Foggy Dataset Compared to MSCOCO?
There are mainly two reasons: 1. The Foggy dataset is a collection of images captured in foggy weather conditions, which inherently falls under the category of complex scenes (such as nighttime, underwater, rainy conditions, etc.). Due to reduced visibility in foggy environments, many objects in the images become blurred, increasing the difficulty for models to accurately recognize them. 2. The training set of MSCOCO is commonly used for training various LVLMs, while datasets of the Foggy type are rarely used for training. Therefore, we can see Foggy as a out-of-distribution test setting. This results in the various LVLMs being unfamiliar with the Foggy style.
@@ 5 Related Work
Large Vision-Language Model
@@
With the success of pretraining in LLMs Touvron et al. (2023) and VFMs Awais et al. (2023), many researchers Alayrac et al. (2022); Li et al. (2023a) extended LLMs to understand real-world images through LVLMs, benefiting from in-context and few-shot learning. This led to a rise in visual instruction-adapted LVLMs Liu et al. (2023c); Zhu et al. (2023); Dai et al. (2023); Gong et al. (2023), which show strong generalization across VL tasks. Most use GPT-4 to generate multimodal instruction datasets and multi-stage pretraining to align visual data with LLMs. For example, Liu et al. Liu et al. (2023c) aligned LLaMA Touvron et al. (2023) with a visual encoder output, while Zhu et al. Zhu et al. (2023) fine-tuned Vicuna Peng et al. (2023) for cross-modal alignment. Similarly, Multimodal GPT Gong et al. (2023) and InstructBLIP Dai et al. (2023) used VL datasets, with the former using BLIP2 Li et al. (2023b) and the latter starting from Flamingo Alayrac et al. (2022).
Despite these advances, LVLMs still struggle with hallucinations in textual output, limiting their effectiveness in vision-language tasks Rohrbach et al. (2018).
Hallucination in LVLMs
Recent studies have focused on hallucination in LVLMs. Some works, summarized in Table 1, address hallucination detection and evaluation Li et al. (2023d); Wang et al. (2023b, a); Jing et al. (2023), while others propose mitigation methods Liu et al. (2023b); Zhou et al. (2023); Yin et al. (2023); Jing and Du (2024); Jing et al. (2025). For instance, Bingo Cui et al. (2023b) evaluates GPT-4V’s hallucinations with bias and interference, and HallusionBench Guan et al. (2024) diagnoses entangled language hallucinations and visual illusions. AutoHallusion Wu et al. (2024) generates benchmarks by manipulating images to challenge language priors, and Hal-Eval Jiang et al. (2024b) categorizes hallucinations into objects, attributes, relations, and events.
Despite progress, fine-grained detection is less explored. Li et al. Li et al. (2023d) introduced POPE to evaluate object-level hallucinations, showing LVLMs’ susceptibility. Wang et al. Wang et al. (2023b) proposed HaELM, creating a hallucination dataset and fine-tuning LLaMA for detection. These methods focus on object-level issues or require training. To address limitations, Wang et al. Wang et al. (2023a) developed AMBER, a benchmark for generative and discriminative tasks involving object, attribute, and relation hallucinations, though it depends on human annotations.
Existing methods mostly primarily use LLMs to extract key information from image captions, such as objects, colors, positions, etc. Based on the extracted keywords and different prompts, the LLM then generates corresponding questions. For each generated question, the LLM also provides the corresponding answer. It can be seen that every step relies entirely on the LLM and manually designed prompts. In contrast, our method does not rely on LLMs for any step, from key information extraction to question and answer generation. When generating questions, it fills in pre-defined, well-structured templates, ensuring both accuracy and controllable question types. Therefore, our method is effeicent and not costly.
@@ 6 Conclusion
Liminations
FIHA has comprehensive features and maintains a high overall quality. Despite the limitations discussed in the previous analysis section, there are additional constraints in some aspects. The generated Q&A primarily focuses on the existence, attributes, and relations of main objects in the images, while lacking in Q&A for surrounding and minor objects. This is due to the FRCNN’s lower confidence in detecting small and less obvious objects.
@@ Appendix B Example of foggy Cityscapes Images datasets
Appendix C Expalnation for DSG
Depending on the answers, some questions in the hallucination benchmark become invalid and thus should not be asked to the LVLM to evaluate hallucination. As shown in Figure 8, if the answer to “are there any flowers here?” is no, dependent questions like “are the flowers white?” should be skipped – the LVLM may often say “the flower doesn’t exist, but it is white”.
Figure 8: The diagram on the right is a schematic illustration of the impact on the results after the introduction of DSG.